OpenStudy (anonymous):

Use inductive reasoning to find the next term in the sequence 2, 3, 5, 8, 12 … What is the next term? A. 13 B. 15 C. 17 D. 18

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle 2+3=~? }\)

OpenStudy (anonymous):

5

OpenStudy (solomonzelman):

so we once saw, that the next term is the sum of the two terms before it. (that is, the 3rd term 5, is the sum of the 1st term 2 and the 2nd term 3)

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle 3+5=~? }\)

OpenStudy (anonymous):

8

OpenStudy (solomonzelman):

yes, and here we see again the same thing some term = sum of the two terms right before it

OpenStudy (solomonzelman):

are you sure you have written the sequence correctly? Are you sure the last number in sequence is 12 and not 13?

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

oh, then it follows a different patten \(\large\color{black}{ \displaystyle 2+\color{red}{1}=3}\) \(\large\color{black}{ \displaystyle 3+\color{red}{2}=5}\) \(\large\color{black}{ \displaystyle 5+\color{red}{3}=8}\) \(\large\color{black}{ \displaystyle 8+\color{red}{4}=12}\)

OpenStudy (solomonzelman):

see that your difference (that you are adding, that I colored in red) is growing by 1 each time?

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

so what would you add to 12 to get to the next term?

OpenStudy (anonymous):

5

OpenStudy (solomonzelman):

yes, and so your next term after 12 is?

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle 12+\color{red}{5}=~?}\)

OpenStudy (anonymous):

17

OpenStudy (solomonzelman):

yup

OpenStudy (anonymous):

thx

OpenStudy (solomonzelman):

yw